a water concentration of calcium carbonate of 5.34x10^-3 mol/L. convert the concentration into mg/L
Answer #1

concentration of calcium carbonate of 5.34x10^-3 mol/L

molar mass of CaCO3         = 100g/mole

concentration of calcium carbonate in g/L   = 5.34x10^-3 mol/L * 100g/mole

                                                                      = 5.34*10^-1g/L

                                                                     = 5.34*10^-1 * 1000mg/L

                                                                      = 534mg/L >>>>>>answer